1. Jamie wanted to find the approximate mass of a bag of rice using a plank and metal weights. She placed the bag of rice on one of the plank`s end and an empty box on the other end. She balanced the plank by filling the box with 25 pieces of metals weights, each weighing 4.4 lbs. What is the mass in kg of the bag of rice? (ANSWER IN G.R.E.S.A.)


1
Expert's answer
2021-03-01T12:39:05-0500

Given:

Bag of rice was balanced by 25 pieces of metals weights, each weighing 4.4 lbs.

Required:

Mass of bag of rice.

Equation:

"m_{rice}=N_{pieces}\times Weight\ of\ one\ piece"

Solution:

"m_{rice}=25\times 4.4\ lbs\times\dfrac{0.45359\ kg}{1\ lbs}=50\ kg."

Answer:

"m_{rice}=50\ kg."